OpenStudy (anonymous):

[9.02] What are the solutions of 3x2 = 18x?

OpenStudy (anonymous):

3x^2?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Ok, so if it's 3x^2 = 18x (^ means squared), then it's simple factoring. You subtract 18x from that side and make it look like so 3x^2 - 18x = 0 From there you take a greatest common factor, since if you factor like for example this: (x-2)(x-2) or something like that, you would always have a third number in the equation which you don't want. The greatest factor is 3x, the factor that both 3x^2 and 18x has. 3x(x-6) = 0 And then you solve for x. Don't forget that there are multiple solutions to this if it looks like that. 3(0)(x-6) = 0 0(x-6) = 0 [Distributing/multiplying 0 to all in the parenthesis makes it 0) Now for the other x.. (x - 6) = 0 [Notice that the 3x isn't there, it's because once you get 0 in a equation and something is multiplied by it, it's automatically 0] x = 6 Final Answer: x = 0 x = 6

OpenStudy (anonymous):

oh my goodness you angel ! you wrote that all out?!?! Thank you so much !!

OpenStudy (anonymous):

No problem :) I just like being organized and making my voice clear.

OpenStudy (anonymous):

I actually understood it thank you thank you

OpenStudy (anonymous):

Solve x2 + 9x + 20 = 0. x = 2, x = 10 x = −2, x = −10 x = −5, x = −4 x = 5, x = 4

OpenStudy (anonymous):

would apply to this too ? @anisanox1

OpenStudy (anonymous):

Would you like for me to explain that one too?

OpenStudy (anonymous):

A "quadratic" is a polynomial that looks like "ax^2 + bx + c", where "a", "b", and "c" are just numbers. For the easy case of factoring, you will find two numbers that will not only multiply to equal the constant term "c", but also add up to equal "b", the coefficient on the x-term. Site here: http://www.purplemath.com/modules/factquad.htm

OpenStudy (anonymous):

perfect thankyoU!